# supplement

### v1

@@ 4 Related work and discussion
In this section, we discuss the obtained bounds in view of known results from the literature. In particular, in Section 4.1 we detail known bounds that are recovered as special cases of Theorem 1. In Sections 4.2 and 4.3, we compare Theorem 1 with deterministic bounds, obtained for bounded noise sequences, and probabilistic bounds for noise sequences with an assumed probability distribution, respectively.
@@ 4.1 Recovering existing bounds as particular cases
We first elucidate the connection between Assumption 1 on the deterministic noise function and energy-boundedness of the noise sequence. As a straightforward consequence of the Representer Theorem [Wahba, 1990; Schölkopf et al., 2001], for the presented results in this paper, the values of the unknown noise function outside of the training input locations are irrelevant (see Appendix A): there exists a noise-generating function $w^{\mathrm{tr}}$ for the data set $\mathcal{D}$ satisfying Assumption 1 if and only if the minimum-norm interpolant $w^{\mu}$ of the (unknown) noise realizations $w_{\mathbb{X}}\doteq\begin{bmatrix}w(x_{1}),\ldots,w(x_{N})\end{bmatrix}^{\top}$ satisfies the RKHS-norm bound $\|w^{\mu}\|_{\mathcal{H}_{k^{w}}}^{2}<\Gamma_{w}^{2}$, where
@@
Therefore, instead of imposing a maximum RKHS norm on $w^{\mathrm{tr}}$, one could equivalently assume a bounded RKHS norm for the minimum-norm interpolant generating the data set. For the Dirac noise kernel $k^{w}(x,x^{\prime})=\delta(x-x^{\prime})$, since $K^{w}_{1:N,1:N}=I_{N}$, the bounded-RKHS-norm assumption on the noise function implies
$\|w^{\mu}\|_{\mathcal{H}_{k^{w}}}^{2}=\sum_{i=1}^{N}w^{\mathrm{tr}}(x_{i})^{2}<\Gamma_{w}^{2},$ (16)
i.e., bounded energy of the noise sequence.
The assumption of bounded energy for the data set has been employed by Fogel [1979] to obtain bounds for the latent function in the setting of linear regression, i.e., finite-dimensional hypothesis spaces. Using the notation adopted in Section 3.3, the non-falsified parameter set is obtained (see [Fogel, 1979, Eq. (3)]) as
$\|\theta^{\mathrm{tr}}-\theta^{\mu}\|_{P^{-1}}^{2}$ $\leq\Gamma_{w}^{2}-y^{\top}y+\|\theta^{\mu}\|_{P^{-1}}^{2}.$ (17)
Proposition 2 shows that the obtained bound recovers this known result from set-membership estimation for finite-dimensional hypothesis spaces. In fact, for the Dirac noise kernel, the worst-case realization of the unknown parameters is given by $\theta^{\star}$ in (12), for which (17) holds with equality. Note that the optimal bound in Theorem 1 does not only recover the bounds by Fogel [1979] for the linear-regression case, but, additionally, provides bounds under the additional complexity constraint $\|\theta^{\mathrm{tr}}\|_{2}^{2}\leq\Gamma_{f}^{2}$.
@@
Building upon the kernel interpolation bound by Weinberger and Golomb [1959] (see [Fasshauer and McCourt, 2015, Section 9.3]), under Assumption 1 the following bound can be derived [Maddalena et al., 2021, Proposition 1]:
@@
where $f^{\mu}_{0}$ is the minimum-norm interpolant in the RKHS $\mathcal{H}_{k^{f}}$ (see (6)) and $\sqrt{\Sigma^{f}_{0}(x_{N+1})}$ is commonly referred to as the power function. The relaxed bound in Lemma 1 generalizes this result: for noise-free measurements, i.e., $\Gamma_{w}^{2}=0$, and in the limit for the noise parameter $\sigma\rightarrow 0$, (18) is recovered exactly by noting that the bound in Lemma 1 is symmetric around the estimate.
4.2 Deterministic bounds
@@
For the Dirac noise kernel $k^{w}(x,x^{\prime})=\delta(x-x^{\prime})$, uncertainty bounds have also been obtained by [Kanagawa et al., 2018, Section 3.4] based on the minimum-norm interpolant $g^{\mu}_{\sigma}\in\mathcal{H}_{k^{f}+k^{\sigma}}$ using the sum of kernels $k^{f}+k^{\sigma}$. Utilizing the fact that, for all $x\notin\mathbb{X}$, the GP posterior mean (5a) and covariance (5b) are equal to the interpolant $g^{\mu}_{\sigma}$ and the corresponding power function, respectively, a bound on the true data-generating function $g\in\mathcal{H}_{k^{f}+k^{\sigma}}$ has been established by [Kanagawa et al., 2018, Proposition 3.8]. However, the bound does not take into account the actual value of the measurements $\{g(x_{i})=y_{i}\}_{i=1}^{N}$, but rather the worst-case realization thereof, rendering it conservative. Additionally, the bound is only valid for the data-generating process $g$ and does not provide bounds for the latent function.
@@
As energy-boundedness is a weaker assumption than point-wise boundedness of the noise, Theorem 1 can also be applied in the setting of point-wise bounded noise, see Section 4.1. In this setting, [Maddalena et al., 2021; Reed et al., 2025] provide closed-form, yet conservative, bounds for the latent function under an RKHS-norm constraint on the latter. The bounds are improved upon by Scharnhorst et al. [2023], which provides optimal point-wise bounds for the latent function. As the bounded-energy and pointwise-boundedness assumptions are equivalent for $N=1$ data points, so are the bounds by Scharnhorst et al. [2023] and Theorem 1 in this case. For larger data sets, under the point-wise-boundedness assumption, the optimal bounds by Scharnhorst et al. [2023] are tighter than the optimal bounds in Theorem 1 obtained under the weaker bounded-energy assumption. Still, their computation relies on solving a constrained convex program, cf. [Scharnhorst et al., 2023, Eq. (6)], whose number of optimization variables is proportional to the number of training data points $N$. Additionally. this optimization problem has to be solved to optimality in order to obtain valid bounds for the latent function, while optimization over the noise parameter $\sigma$ in (9) returns a valid upper bound for all $\sigma\in(0,\infty)$.
4.3 Probabilistic bounds
@@
Compared to Theorem 1, these bounds hold with a user-chosen probability $p\in(0,1)$, use a fixed constant $\sigma>0$, but otherwise have the same structure and the same assumption on the latent function $f^{\mathrm{tr}}$ in terms of a known bound on its RKHS norm. However, while the proposed analysis considers energy bounded noise, $\|w^{\mathrm{tr}}\|_{\mathcal{H}_{k^{w}}}^{2}<\Gamma_{w}^{2}$ (Assumption 1), these results apply to (conditionally) independent sub-Gaussian noise [Srinivas et al., 2012; Abbasi-Yadkori, 2013; Fiedler et al., 2021]. This is a stronger [^1] To be precise, if the noise is sub-Gaussian, we can derive an energy bound $\Gamma_{w}^{2}$, such that $w^{\mathrm{tr}}$ satisfies Assumption 1 with the kernel $k^{w}(x,x^{\prime})=\delta(x-x^{\prime})$ and a desired probability $p\in(0,1)$. However, the converse is not true as energy-bounded noise may be correlated and biased.requirement as it does not allow for biased or correlated noise, which can be difficult to ensure in real-world experiments. Nonetheless, both the proposed bound and existing high-probability bounds can be applied in case of independent, zero-mean and bounded noise; in the following, we numerically investigate the conservativeness of the bounds in this setting.
@@
Using a squared-exponential kernel $k^{f}(x,x^{\prime})=\exp(-\|x-x^{\prime}\|^{2})$ for the latent function, as well as a Dirac noise kernel $k^{w}(x,x^{\prime})=\delta(x-x^{\prime})$ on the domain $\mathcal{X}=[0,4]$, random latent functions are generated with $\|f^{\mathrm{tr}}\|_{\mathcal{H}_{k^{f}}}^{2}=\Gamma_{f}^{2}=1$. Training data is sampled based on measurement noise following a zero-mean truncated Gaussian distribution with standard deviation and bounded absolute value equal to $\epsilon=0.01$, which is $R$-sub-Gaussian for $R=\epsilon$. The corresponding bound for the noise energy is derived as $\Gamma_{w}^{2}=N\epsilon^{2}$. We compare the proposed bound (Theorem 1), which is optimal given only the information $\|w^{\mathrm{tr}}\|_{\mathcal{H}_{k^{w}}}^{2}\leq\Gamma_{w}^{2}$, the relaxed bound (Lemma 1) with $\sigma=\epsilon$, and a standard high-probability error bound Abbasi-Yadkori [2013], cf. [Fiedler et al., 2024, Eq. (7)], that uses only sub-Gaussianity of the noise and provides a valid bound with probability $p=1-\delta=99\%$, similar to [Srinivas et al., 2012; Fiedler et al., 2021]. Figure 2 compares the area of the confidence region for a varying number of randomly sampled training points $N=1,\dots,10^{3}$, averaged over 1000 runs for randomly sampled unknown functions $f^{\mathrm{tr}}$. To guarantee optimality of the optimal upper and lower bounds in Theorem 1, they are determined using convex, finite-dimensional reformulations of (2), (3), cf. Appendix A; the optimization problems are solved using CVXPY [Diamond and Boyd, 2016; Agrawal et al., 2018]. The relaxed and probabilistic bounds respectively require evaluation of Eqs. 5, 6 and 7. All runs are performed in parallel on a 14-core Intel i9-7940X CPU, with an overall runtime of about 18h for the entire experiment. In the low data regime, the proposed optimal and relaxed bounds, leveraging energy-boundedness of the noise, are significantly less conservative compared to the probabilistic bound. However, as multiple similar data points may provide no additional information without probabilistic information, they do not significantly improve after a certain number of data points $N$. In contrast, the probabilistic bound leverages independence and hence asymptotically attains smaller uncertainty bounds with increasing data. However, it should be noted that these probabilistic bounds are only valid if indeed the noise is (conditionally) independent and zero mean and otherwise these shrinking confidence intervals may be misleading.
@@
Figure 2: Numerical comparison of area of uncertainty region for increasing number of data points $N$ with $\{5\%,95\%\}$-percentiles shown in shade.

### v3

@@ 4 Related work and discussion
In this section, we discuss the obtained bounds also in view of known results from the literature. In particular, in Section˜4.1 we detail known bounds that are recovered as special cases of Theorem˜1. In Sections˜4.2 and 4.3, we respectively compare Theorem˜1 with deterministic bounds, obtained for bounded noise sequences, and probabilistic bounds, for noise sequences with an assumed probability distribution. Specifically, Section˜4.3 encompasses a thorough numerical comparison of the bounds, including an application on safe control that shows the effectiveness of the proposed bounds on a downstream task.
@@ 4.1 Recovering existing bounds as particular cases
We first elucidate the connection between Assumption 1 on the deterministic noise function and energy-boundedness of the noise sequence. As a straightforward consequence of the Representer Theorem [Wahba, 1990; Schölkopf et al., 2001], for the presented results in this paper, the values of the noise function outside the training input locations are irrelevant (see Appendix˜A): there exists a noise-generating function $w^{\mathrm{tr}}$ for the data set $\mathcal{D}$ satisfying Assumption˜1 if and only if the minimum-norm interpolant $w^{\mu}$ of the (unknown) noise realizations $w_{\mathbb{X}}\doteq\begin{bmatrix}w(x_{1}),\ldots,w(x_{N})\end{bmatrix}^{\top}$ satisfies the RKHS-norm bound $\|w^{\mu}\|_{\mathcal{H}_{k^{w}}}^{2}<\Gamma_{w}^{2}$, where
@@
Therefore, instead of imposing a maximum RKHS norm on $w^{\mathrm{tr}}$, one could equivalently assume a bounded RKHS norm for the minimum-norm interpolant generating the data set. For the Dirac noise kernel $k^{w}(x,x^{\prime})=\delta(x-x^{\prime})$, since $K^{w}_{1:N,1:N}=I_{N}$, the bounded-RKHS-norm assumption on the noise function implies bounded energy of the noise sequence, i.e.,
$\|w^{\mu}\|_{\mathcal{H}_{k^{w}}}^{2}=\sum_{i=1}^{N}w^{\mathrm{tr}}(x_{i})^{2}<\Gamma_{w}^{2}.$
The assumption of bounded energy for the data set has been employed by Fogel [1979] to obtain bounds for the latent function in the setting of linear regression, i.e., finite-dimensional hypothesis spaces. Using the notation adopted in Section˜3.3, the non-falsified parameter set is obtained as
$\|\theta^{\mathrm{tr}}-\theta^{\mu}\|_{P^{-1}}^{2}\leq\Gamma_{w}^{2}-y^{\top}y+\|\theta^{\mu}\|_{P^{-1}}^{2},$ (16)
see [Fogel, 1979, Eq. (3)]. Proposition˜2 shows that the obtained bound recovers this known result from set-membership estimation for finite-dimensional hypothesis spaces. In fact, for the Dirac noise kernel, the worst-case realization of the unknown parameters is given by $\theta^{\star}$ in (12), for which (16) holds with equality. Note that the optimal bound in Theorem˜1 does not only recover the bounds by Fogel [1979] for the linear-regression case, but, moreover, provides bounds under the additional complexity constraint $\|\theta^{\mathrm{tr}}\|_{2}^{2}\leq\Gamma_{f}^{2}$.
@@
Building upon the kernel interpolation bound by Weinberger and Golomb [1959] (see [Wendland, 2004, p. 192][Fasshauer and McCourt, 2015, Section 9.3]), under Assumption 1 the following bound can be derived, cf. [Maddalena et al., 2021, Proposition 1]:
@@
where $f^{\mu}_{0}$ is the minimum-norm interpolant in the RKHS $\mathcal{H}_{k^{f}}$ (see (6)) and $\sqrt{\Sigma^{f}_{0}(x_{N+1})}$ is commonly referred to as the power function. The relaxed bound in Lemma˜1 generalizes this result: for noise-free measurements, i.e., $\Gamma_{w}^{2}\rightarrow 0$, and in the limit for the noise parameter $\sigma\rightarrow 0$, (17) is recovered exactly by noting that the bound in Lemma˜1 is symmetric around the estimate.
4.2 Comparison with existing deterministic bounds
@@
For the Dirac noise kernel $k^{w}(x,x^{\prime})=\delta(x-x^{\prime})$, uncertainty bounds have also been obtained by [Kanagawa et al., 2025, Section 6.4] based on the minimum-norm interpolant $g^{\mu}_{\sigma}\in\mathcal{H}_{k^{f}+k^{\sigma}}$ using the sum of kernels $k^{f}+k^{\sigma}$. Utilizing the fact that, for all $x\notin\mathbb{X}$, the GP posterior mean (5a) and covariance (5b) are equal to the interpolant $g^{\mu}_{\sigma}$ and the corresponding power function, respectively, a bound on the true data-generating function $g\in\mathcal{H}_{k^{f}+k^{\sigma}}$ has been established by [Kanagawa et al., 2025, Corollary 6.8]. However, the bound does not take into account the actual value of the measurements $\{g(x_{i})=y_{i}\}_{i=1}^{N}$, but rather the worst-case realization thereof, rendering it conservative. Additionally, the bound is only valid for the data-generating process $g$ and does not provide bounds for the latent function.
@@
As energy-boundedness is a weaker assumption than point-wise boundedness of the noise, Theorem˜1 can also be applied in the setting of point-wise bounded noise, see Section˜4.1. In this setting, [Maddalena et al., 2021; Reed et al., 2025] provide closed-form, yet conservative, bounds for the latent function under an RKHS-norm constraint on the latter. The bounds are improved upon by Scharnhorst et al. [2023], which provides optimal point-wise bounds for the latent function. As the bounded-energy and pointwise-boundedness assumptions are equivalent for $N=1$ data points, so are the bounds by Scharnhorst et al. [2023] and Theorem˜1 in this case. For larger data sets, under the point-wise-boundedness assumption, the optimal bounds by Scharnhorst et al. [2023] are tighter than the optimal bounds in Theorem˜1 obtained under the weaker bounded-energy assumption. Still, their computation relies on solving a constrained convex program, cf. [Scharnhorst et al., 2023, Eq. (6)], whose number of optimization variables is proportional to the number of training data points $N$. Additionally. this optimization problem has to be solved to optimality in order to obtain valid bounds for the latent function, while optimization over the noise parameter $\sigma$ in (9) returns a valid upper bound for all $\sigma\in(0,\infty)$.
4.3 Comparison with existing probabilistic bounds
@@
Compared to Theorem 1, these bounds hold with a user-chosen probability $p\in(0,1)$, use a fixed constant $\sigma>0$, but otherwise have the same structure and the same assumption on the latent function $f^{\mathrm{tr}}$ in terms of a known bound on its RKHS norm. However, while the proposed analysis considers energy-bounded noise, $\|w^{\mathrm{tr}}\|_{\mathcal{H}_{k^{w}}}^{2}<\Gamma_{w}^{2}$ (Assumption 1), these results apply to (conditionally) independent sub-Gaussian noise [Srinivas et al., 2012; Abbasi-Yadkori, 2013; Fiedler et al., 2021]. This is a stronger [^1] To be precise, if the noise is sub-Gaussian, we can derive an energy bound $\Gamma_{w}^{2}$, such that $w^{\mathrm{tr}}$ satisfies Assumption 1 with the kernel $k^{w}(x,x^{\prime})=\delta(x-x^{\prime})$ and a desired probability $p\in(0,1)$. However, the converse is not true as energy-bounded noise may be correlated and biased. requirement as it does not allow for biased or correlated noise, which can be difficult to ensure in real-world experiments. Nonetheless, both the proposed bound and existing high-probability bounds can be applied in case of independent, zero-mean and bounded noise; in the following, we numerically investigate the conservativeness of the bounds in this setting [^2] The code to reproduce the experiments is publicly available at https://gitlab.ethz.ch/ics/bounded-energy-rkhs-bounds and at https://doi.org/10.3929/ethz-c-000785083..
@@
In this experiment, the size of the uncertainty regions is compared. Using a squared-exponential kernel $k^{f}(x,x^{\prime})=\exp(-\|x-x^{\prime}\|^{2}/\ell^{2})$, $\ell=1$, for the latent function, as well as a Dirac noise kernel $k^{w}(x,x^{\prime})=\delta(x-x^{\prime})$ on the domain $\mathcal{X}=[0,4]$, random latent functions are generated with $\|f^{\mathrm{tr}}\|_{\mathcal{H}_{k^{f}}}^{2}=\Gamma_{f}^{2}=1$. Training data are sampled based on measurement noise following a zero-mean truncated Gaussian distribution with standard deviation and bounded absolute value equal to $\epsilon=0.01$, which is $R$-sub-Gaussian for $R=\epsilon$. The corresponding noise-energy bound is derived as $\Gamma_{w}^{2}=N\epsilon^{2}$. We compare the proposed bound (Theorem 1), which is optimal given only the information $\|w^{\mathrm{tr}}\|_{\mathcal{H}_{k^{w}}}^{2}\leq\Gamma_{w}^{2}$, the relaxed bound (Lemma 1) with $\sigma=\epsilon$, and a standard high-probability error bound [Abbasi-Yadkori, 2013], cf. [Fiedler et al., 2024, Eq. (7)], which uses only sub-Gaussianity of the noise and provides a valid bound with probability $p=0.99$, similar to [Srinivas et al., 2012; Fiedler et al., 2021]. Optimality of the numerical solution to (9) is guaranteed by solving a convex reformulation of (2) (see Appendix˜A) using CVXPY [Diamond and Boyd, 2016; Agrawal et al., 2018]. Figure 2 compares the area of the uncertainty region for $N=1,\dots,10^{3}$ randomly sampled training points, averaged over $10^{3}$ runs for randomly sampled latent functions $f^{\mathrm{tr}}$. In the low-data regime, the proposed optimal and relaxed bounds, leveraging energy-boundedness of the noise, are significantly less conservative. However, as multiple similar data points may provide no additional information without probabilistic information, they do not significantly improve after a certain number of data points $N$. In contrast, the probabilistic bound leverages independence, asymptotically attaining smaller uncertainty bounds with increasing data. However, it should be noted that these probabilistic bounds are only valid if indeed the noise is (conditionally) independent and zero-mean; otherwise, these shrinking confidence intervals would be misleading.
@@
Figure 2: Numerical comparison of area of uncertainty region for increasing number of data points NN with $\{5\%,95\%\}$-percentiles shown in shade.
Safe control for uncertain nonlinear systems
Lastly, we demonstrate the application of the proposed bounds to the downstream task of safe control. Consider the uncertain (for simplicity scalar) nonlinear dynamical system
$x(k+1)$ $=f^{\mathrm{known}}(x(k),u(k))+f^{\mathrm{tr}}(x(k),u(k)),$ (18)
with known dynamics $f^{\mathrm{known}}$ and unknown residual dynamics $f^{\mathrm{tr}}$. Given the current state $x(k)\in\mathbb{R}$ of the system at time $k\in\mathbb{N}$, the goal is to find an optimal control input $u(k)\in\mathbb{R}$ that minimizes a user-defined cost function $c(x(k),u(k))$, subject to a safety-critical constraint, $f^{\mathrm{known}}(x(k),u(k))+f^{\mathrm{tr}}(x(k),u(k))\geq(1-\gamma)x(k)$ – similar to a control barrier function [Agrawal and Sreenath, 2017; Ames et al., 2019; Jagtap et al., 2020]. The uncertainty in the system dynamics is handled by leveraging the proposed (the probabilistic) bound to enforce robust constraint satisfaction for all functions in the uncertainty set, containing the unknown function $f^{\mathrm{tr}}$ (with probability $p$). Importantly, this makes tight uncertainty bounds desirable, since they generally lead to lower costs and a larger feasible region, where safety of the control input can be guaranteed.
We compare the bounds for the following example setup: $f^{\mathrm{known}}(x,u)\doteq 0.5x+u-1$, $f^{\mathrm{tr}}(x,u)\doteq\exp(-x^{2})\sin(10x)$, $c(x,u)\doteq(f^{\mathrm{known}}(x,u)+f^{\mu}_{\sigma}(x,u))^{2}+u^{2}$, $k^{f}$, $k^{w}$ as above with $\ell=\sqrt{2}/20$. For the proposed bound, the optimization problem is formulated using the relaxed bound in Eq.˜8, optimizing $\sigma\in(0,\infty)$ and $u(k)\in[-2,2]$ simultaneously; for the probabilistic bounds, $\sigma=\epsilon$ is fixed with the same noise assumptions as in Section˜4.3; see Appendix˜D for implementation details. Additionally, we implement the proposed bound using only the nearest 10 training points to construct the uncertainty bounds. Fig.˜3 shows the success rate in terms of the share of feasible problems for an increasing amount of training data on a grid of $500$ test points in the domain $x(k)\in[-2,2]$, repeated $20$ times with random noise realizations. Due to the smaller uncertainty bounds, the proposed bound using the full data set achieves the highest success rate, albeit at a high computational cost. In contrast, using the probabilistic bounds and the subset-of-data variant of the proposed bounds leads to similar success rates, with the latter exhibiting significantly lower computation times due to its independence of the number of training points. Note that utilizing the probabilistic bounds with test-point-dependent subsets of data would generally deteriorate the probability $p$ of their joint validity for all test points, compromising the controller’s safety guarantees.
Figure 3: Application of uncertainty bounds for safe control. Success rate (upper plot) and solve time (lower plot) with $\{5\%,95\%\}$-percentiles shown in shade.
4.4 Limitations
The obtained bounds suffer from common criticalities and limitations of kernel-based learning, which are (a) dealing with kernel mis-specification and (b) knowing valid RKHS-norm bounds. Both issues are typically addressed empirically, (a) by hyper-parameter tuning via cross-validation or maximum-likelihood estimation ([Wahba, 1990], [Rasmussen and Williams, 2006, Section 5.4], [Karvonen et al., 2020]) and (b) by estimating the bound value from data [Csáji and Horváth, 2022; Tokmak et al., 2024]. While a rigorous investigation of (a) is beyond the scope of this paper, we note that mis-specification of $k^{f}$ may also be compensated by an inflated RKHS-norm bound $\Gamma_{w}$. Regarding (b), for the latent function, this is a common assumption in the literature on kernel-based uncertainty bounds; for the noise function, we discuss in Section˜4.1 how it generalizes the common setting of energy-bounded noise. Under-estimation in the RKHS-norm bounds could be detected by checking feasibility of the optimization problem. Conversely, we point out that considering conservative values of $\Gamma_{f}$ and $\Gamma_{w}$ merely results in a sublinear inflation of the computed uncertainty envelope, cf. Eq.˜7. Nevertheless, further research will be devoted to rigorously assessing the robustness of the obtained bounds with respect to possible mis-specifications in (a) and (b).